OpenStudy (anonymous):

What is the equation of the line passing through (8, –6) and perpendicular to the line y = x – 2 in slope-intercept form?

OpenStudy (anonymous):

the slope of given line is m=1 so slope of required perpendicular line will be m=-1 use point slope form y-y0=m(x-x0) y-(-6)=-1(x-8) y+6=-x+8 y=-x+8-6 y=-x+2